### v1

# A non-Hermitian loop for a quantum measurement

Abstract
We establish a framework for a mechanism steering state vector collapse through time evolution as described by the Schrödinger’s equation. Under the premise of a non-Hermitian Hamiltonian dynamics, a main outcome is that collapse has to occur when the Hamiltonian completes a closed loop in the parameter space encoding the interaction with the meter. For two-level systems, we put forward the phenomenon of chiral state conversion as a mechanism effectively eliminating superpositions. Our results open a way to simulate quantum measurements in classical systems that up to now were restricted to the Schrödinger part of the quantum dynamics.
Introduction and overview of key findings.– One century after its inception, quantum mechanics stands as one of humanity’s most rigorously tested theories, yet it still keeps a foundational challenge known as the measurement problem Bell (1990); Dirac (1963); Laloë (2001). The crux of the measurement problem lies in explaining or making sense of the wave function collapse. The postulates of quantum mechanics state that during a measurement, the Schrödinger equation, which ordinarily governs quantum dynamics, is momentarily suspended to allow the wave function to collapse (see for example Cohen-Tannoudji et al. (2020); Shankar (1994)). This split description of the quantum dynamics is introduced to bridge the complex world of quantum superpositions with the familiar classical reality, where those superpositions have vanished. The incompatibility between the two descriptions for the dynamics entails a “cut” between the quantum and classical realms (as Werner Heisenberg called it in a posthumously published letter Bacciagaluppi and Crull (2011) in response to the EPR paper Einstein et al. (1935)). Indeed, the collapse is essentially a non-unitary, irreversible operation with loss of information and eventual energy and momentum exchange Pearle (2000), or as John Wheeler put it “an irreversible act of amplification such as the blackening of a grain of silver bromide emulsion” Wheeler (2016)  [^1] Wheeler mentions that the term irreversible amplification was used by Bohr, see chapter 1 of Bohr (2010), which was published in 1958..
Despite advances in understanding the quantum-classical transition Zurek (2003); Schlosshauer (2007), the measurement problem remains a mystery for many of us Leggett (2006); Hance and Hossenfelder (2022)  [^2] There are also notable opinions against thinking of the measurement problem as such, see for example Mermin (2022). We also refer to Refs. Schlosshauer (2011); Schlosshauer et al. (2013) for a broader overview on the different positions.. Two proposals to address the conventional view of measurements are the Ghirardi-Rimini-Weber (GRW) Ghirardi et al. (1986, 1990) and the Continuous Spontaneous Localization (CSL) models Pearle (1989). Both introduce stochastic elements to the Schrödinger’s equation, suggesting that wave function collapse can occur spontaneously Bassi and Ghirardi (2003) (there are also other proposals using stochastic elements such as Fernández-Alcázar and Pastawski (2015)). Although GRW models do not rely on the concept of an apparatus to induce collapse, they still lack a microscopic explanation for how collapse occurs. Other proposals, inspired by GRW, are due to Diosi Diósi (1989); Donadi et al. (2021) and Penrose Penrose (1996) linking the wave function collapse to gravity. More recently, different authors (see  Martínez Romeral et al. (2024); Singh et al. (2023a, b) and also Van Wezel (2010); Mertens et al. (2023); Mukherjee and van Wezel (2024)) proposed using a non-Hermitian Hamiltonian to model a quantum measurement (earlier proposals also include dissipative modifications of the Schrödinger equation Gisin and Piron (1981)). These models are steps to make sense of quantum collapse Bell (1990), yet they also underscore its complexity Bassi et al. (2013) and the need for further experiments.
In a parallel and unfolding story, over recent years, we have seen a rise in research fronts using photonics Rechtsman et al. (2013); Vicencio et al. (2015); Guzmán-Silva et al. (2014); Stützer et al. (2018), metamaterials Veenstra et al. (2024), acoustics Zhang et al. (2021), mechanical systems Lorenz et al. (2023) and circuits Lee et al. (2018); Imhof et al. (2018) as classical analogs for quantum systems. Notably, in photonic lattices, the paraxial Helmholtz equation for the electromagnetic waves maps to the Schrödinger equation Szameit and Nolte (2010) thus allowing to simulate quantum systems. Similarly, in electrical circuits, the correspondence between the circuit Laplacian and the tight-binding Hamiltonian have been extensively used to investigate topological states Lee et al. (2018); Imhof et al. (2018). Moreover, it has been demonstrated that these circuits can also mimic Schrödinger dynamics Ezawa (2019). Collectively, these examples illustrate the availability of a classical toolkit for quantum system simulation. However, a noteworthy limitation remains apparent: classical simulators have yet to fully address quantum measurements, which, because of the split description mentioned earlier, are distinct from quantum dynamics and thus remain a significant challenge.
In this Letter, we present a theoretical framework for non-Hermitian models of wave function collapse. Non-Hermitian Hamiltonians have been shown to exhibit many interesting features Moiseyev (2011); Longhi (2009); Rotter and Bird (2015). One of them is the appearance of non-Hermitian degeneracies, called exceptional points Kato (1995); Moiseyev (2011), where both eigenvalues and eigenvectors coalesce. Exceptional points are at the heart of many counterintuitive Heiss and Harney (2001); Dembowski et al. (2001); Heiss (2012) and useful phenomena, such as the typical amplification of small parametric variations found near them which may lead to novel sensors  Wiersig (2014); Rechtsman (2017); Hodaei et al. (2017); Chen et al. (2017); Hokmabadi et al. (2019). Furthermore, non-Hermitian Hamiltonians offer a multitude of testing grounds in metamaterials Veenstra et al. (2024), photonics Szameit et al. (2011); Regensburger et al. (2012); Pan et al. (2018); Miri and Alù (2019); Vicencio et al. (2024), acoustics Zhang et al. (2021) and circuits Schindler et al. (2011); Hofmann et al. (2020), thereby offering a unique opportunity of simulating and discussing quantum collapse outside of the purely quantum domain.
To set the stage, we first revisit the postulates of quantum mechanics, laying the groundwork for our discussion. Subsequently, we discuss a set of physically motivated assumptions and their consequences. The main outcome of this part is that, if a quantum measurement is to be modeled as a process that occurs in time and which is described by the Schrödinger equation with a non-Hermitian Hamiltonian encoding the effect of the meter (or whatever interaction leading to collapse), then the Hamiltonian should describe a closed loop in parameter space. Interestingly, we are able to readily link this discovery with the phenomenon called state exchange or chiral state conversion: For the $2\times 2$ case, when moving along a closed contour in the space of the Hamiltonian parameters, the initial state undergoes a conversion to a preferred state, irrespective of initial conditions, which is determined by the sense of rotation in the space of parameters Uzdin et al. (2011); Berry and Uzdin (2011); Hassan et al. (2017a). Thus, chiral state conversion serves the purpose of breaking the superpositions and modeling quantum collapse under a single dynamic equation. However, there is a twist to consider: the sense of the loop is a property of the interaction itself, rather than the initial Hermitian Hamiltonian. In other words, it is an additional variable, and compatibility with standard quantum mechanics requires that its value must be set to satisfy Born’s rule. We close by discussing the implications of our results in the path to simulating a quantum measurement in classical systems.
Split procedures for the quantum dynamics.– Quantum mechanics is traditionally based on four key postulates, as outlined in textbooks Cohen-Tannoudji et al. (2020); Shankar (1994). Two of these postulates specifically address the theory’s dynamics. One of them pertains to the standard, measurement-free behavior governed by the Schrödinger equation, highlighting how systems evolve when not being measured. The other postulate introduces the concept of measurement as a distinct phenomenon, where the act of measuring disrupts the usual dynamics, causing the system’s state vector to collapse into one of the observable’s eigenstates. This collapse results in the system adopting a state that corresponds to one of the observable’s eigenvalues, rendering the measurement as an exceptional process within quantum mechanics, described by a different antagonistic procedure.
Figure 1: (a) Scheme showing a loop in a multidimensional parameter space defining the Hamiltonian. ${\cal H}_{0}$ is the initial Hamiltonian just before the interaction that leads to the collapse starts. (b) Time-evolution of the spin-up (blue) and spin-down (orange) relative probabilities for a system evolving under a Hamiltonian proportional to $\sigma_{x}$. A non-Hermitian term adding gains on the spin-up and loss on the spin-down is turned on with a gaussian function centered at time $t=5$ (standard deviation is $1$). This pulsed perturbation leads to a collapse after which the time-evolution restarts. (c) Scheme showing the Riemann surfaces of the real and imaginary part of the eigenvalues near an exceptional point. (d) Because of the peculiar shape of the solutions in presence of an exceptional point, the phenomenon of chiral state conversion leads to the asymmetric conversion of the initial eigenstates |$\ket{\psi_{1}}$ and |$\ket{\psi_{2}}$ after a closed loop which defies the adiabatic intuition and destroys superposition states.
Assumptions for a dynamical model of a quantum measurement.– Our discussion is based on the following set of assumptions for a family of models based on non-Hermitian Hamiltonians:
  1. 1.
    The measurement takes place over a finite time and the time evolution is dictated by the Schrödinger equation, with an eventual renormalization to keep the norm during the measurement process.
  2. 2.
    Over the timescale of the measurement, the Hamiltonian describes a closed loop in parameter space $\mathbf{X}(t)=\{X_{1}(t),X_{2}(t),...\}$, so that it remains unchanged at the end of the process (Fig. 1 (a)):
    $\begin{split}{\cal H}(\mathbf{X}(t=T))={\cal H}(\mathbf{X}(t=0))={\cal H}_{0}.\end{split}$ (1)
  3. 3.
    The interaction between the system and the measuring apparatus is modeled by introducing a non-Hermitian term to the Hamiltonian. For consistency with the previous points, this term needs to be time-dependent and describe a loop in parameter space. Details of the loop may also depend on the initial state.
Assumption 1 is the most natural one in order to recover unity in the description of the dynamics. Assumption 2 seems a reasonable assumption as we want for the system to remain the same when the apparatus (or the source of interaction) is removed. As a byproduct, the idea of making a closed loop in parameter space is naturally introduced.
Regarding assumption 3, the non-Hermitian term is a minimal model to account for the action of the measurement apparatus on the system without recurring to a larger Hilbert space, probably containing the degrees of freedom of the apparatus and/or the more complicated many-body interactions that escape the simple single-particle picture Foa Torres (2019). Another, more pragmatic motivation is that non-Hermitian Hamiltonians are readily available in quantum simulators using photonic lattices Szameit et al. (2011); Regensburger et al. (2012); Pan et al. (2018); Miri and Alù (2019) and electrical circuits Schindler et al. (2011); Hofmann et al. (2020).
All these assumptions can be summarized in two sentences: The measurement takes place over a finite time and is modeled as a non-Hermitian perturbation to an otherwise Hermitian Hamiltonian ${\mathcal{H}}_{0}$ which describes a loop in parameter space, thereby returning to the initial value after the interaction is complete. The parameter space is composed of those variables defining the Hamiltonian (including those related to the meter) and some of these parameters may have values set by the initial state.
The case of Ref. Martínez Romeral et al. (2024), where the collapse is obtained in a Stern-Gerlach setup (including spin and position variables), does not include an explicit time-dependence. However it can be thought as a particular equivalent case since its role is taken by the passage of the particle through the region where the magnetic field is acting.
Possible mechanisms.– Let us now turn to the discussion of the possible mechanisms. There are different ways where a non-Hermitian term could lead to a collapse of a superposition state on one of its components. A very simple way would be just adding gains and losses as to amplify one of the possible states and reduce the others. A typical case is shown in Fig.1 (b), where a superposition state is collapsed onto one of its components. In this case, we take for illustration purposes:
${\cal H}={\cal H}_{0}+\lambda{\cal H}_{1}.$ (2)
where:
${\cal H}_{0}=\gamma_{0}\begin{pmatrix}0&1\\ 1&0\end{pmatrix},$ (3)
${\cal H}_{1}=\begin{pmatrix}i\Gamma&0\\ 0&-i\Gamma\end{pmatrix}.$ (4)
The time dependence is embedded in the parameter $\lambda$ which is taken as a normalized Gaussian centered at the measurement time ($t=5$ in units of $\hbar/\gamma_{0}$) and with a standard deviation $\sigma=1$. Fig.1 (b) shows the probabilities of having spin up or down (dashed and full lines respectively) as a function of time. The meter acts around $t=5$ amplifying the spin down and reducing the spin up components leading to a situation where the superposition vanishes, the collapse. After the interaction with the meter vanishes the usual time-evolution leads again to a superposition state.
While this scheme may be adjusted to obtain a desired outcome it leaves, however, no much insight on the process of the measurement itself nor the ultimate mechanism leading to the collapse. One may also argue that substantial fine tuning of the non-Hermitian Hamiltonian is necessary to mimic a quantum collapse  [^3] Note that in this case the Hamiltonian describes a loop in parameter space enclosing a vanishing area..
For simplicity, in what follows we will focus on energy measurements A hint for the search for other possible mechanisms, is that the underlying phenomenon needs to bear an intrinsic asymmetry so that by changing a variable in the “measurement apparatus” one could obtain the different possible outcomes, it also needs to be a destroyer of superpositions. Furthermore, the mathematical substrate for such phenomenon is likely to break many of the usual rules to serve as a proper backdrop for the cut. Within the realm of non-Hermitian systems, one of the most prominent sources of unusual and counter-intuitive physics are the so called exceptional points Kato (1995); Heiss (2012). These are points in parameter space where the eigenvalues and eigenvectors coalesce (see the scheme in Fig.1 (c) illustrating the Riemann surfaces close to an exceptional point), and have been the focus of intense research  Dembowski et al. (2001, 2004); Stehmann et al. (2004); Hu et al. (2017); Miri and Alù (2019). The coalescence at an exceptional point needs to be distinguished from a usual Hermitian degeneracy where a full set of eigenstates is kept. The key difference being the appearance of defectiveness Foa Torres (2019), the lack of a full set of eigenstates. The dependence of the eigenvalues on the parameters around exceptional points is typically non-analytic Heiss (2012).
Although for a $2\times 2$ system non-Hermitian system, the adiabatic solutions are expected to exchange when encircling an exceptional point Heiss and Harney (2001); Dembowski et al. (2001); Heiss (2012), this conversion has been predicted to occur in an asymmetric way Uzdin et al. (2011); Berry and Uzdin (2011), the so called chiral state or mode conversion Hassan et al. (2017a) because of non-adiabatic transitions. This phenomenon manifests in a two-state system during the encirclement of an exceptional point within the parameter space. In the realm of Hermitian systems, the adiabatic theorem posits that if a system is initially prepared in one of its eigenstates and the Hamiltonian parameters are gradually altered along a closed trajectory, the system will revert to its initial state, albeit with an additional geometric phase Berry (1984). In contrast, in non-Hermitian systems, a clockwise encirclement around an exceptional point invariably results in the system settling into one of the eigenstates (deemed the preferred state). On the other hand, an anticlockwise journey consistently leads to the emergence of the second eigenstate, as depicted in Fig.1 (d). This intriguing behavior has been observed across various platforms, including waveguides Doppler et al. (2016), and optomechanical systems Xu et al. (2016). At a fundamental level, this state conversion is associated with the Stokes phenomenon of asymptotics Berry and Uzdin (2011) and the concept of stability loss delay Milburn et al. (2015).
Remarkably, the dynamics of the instantaneous eigenstates’ weights in non-Hermitian systems diverge significantly from those in Hermitian counterparts. The weights can exhibit extreme fluctuations Berry and Uzdin (2011), and transitions can occur even under slow dynamics, challenging the conventional understanding of adiabaticity Doppler et al. (2016); Berry and Uzdin (2011). Ref. Hassan et al. (2017a) has presented analytical and numerical insights on why a preferred state gets amplified in a dynamical setting and has shown that the state conversion is robust to deformations of the contour if the initial and final points are kept Hassan et al. (2017a). Recent findings further reveal that chiral state conversion can occur even without the need to encircle the exceptional point Hassan et al. (2017b); Nasari et al. (2022), thus broadening the phenomenon’s observed scope.
In our present context, chiral state conversion provides a promising mechanism for modeling the state vector collapse in a $2\times 2$ system. For this, the interaction with the meter should lead to exceptional points in parameter space. The loop starts (and ends) with the Hamiltonian being ${\cal H}_{0}$ which is Hermitian. Right at the beginning of the loop, a crucial decision must be made regarding the direction in which the loop will proceed. Switching between clockwise and anticlockwise loops changes the outcome or the preferred state. Upon the loop’s completion, when the interaction ceases and the Hamiltonian reverts to its original form ${\cal H}_{0}$, the initial superpositions that may have been in the initial state disappear, ushering in a state of certainty. This transition marks a pivotal moment in the dynamics of the system, where the ephemeral nature of quantum superpositions gives way to definitive outcomes, encapsulating the essence of quantum measurement and evolution.
The probability with which the meter chooses between the clockwise and anticlockwise loops should depend on the initial state and follow Born’s rule if the usual quantum mechanics predictions are to be preserved. This renders the dynamical equation not only non-Hermitian but also non-linear. Another point is that, for a sufficiently slow variation of the parameters, one would not need to fine tune the loop to encircle the exceptional point as the conversion has been shown to occur anyway Hassan et al. (2017b); Nasari et al. (2022).
Collapse when there is no measurement apparatus.– Collapse is known to occur even without a measurement apparatus. When there is a measurement apparatus it seems reasonable to model it as an additional non-Hermitian term that acts during the time of the measurement. But how to account for exceptional points in other cases, as when many body interactions are present? A possible path for an answer could be the proliferation of exceptional points reported in Luitz and Piazza (2019) or the phenomena reported in Žnidarič (2022).
We underline the fact that any interaction that vanishes after some time (this is, that does not modify permanently the Hamiltonian of the system of interest) adheres to assumption 2. This compliance allows for the autonomy of loops within the parameter space, enabling for spontaneous collapse, or collapse without a meter, to occur. In this case, one should just rather view collapse as emerging from the interaction with other system, encoded in an effective way in the loop in parameter space.
Classical simulators for state vector collapse.– As a corollary, and perhaps most immediately useful outcome of our discussion, one can envision the simulation of quantum measurements using classical objects: photonic setups with classical light, electrical circuits and many others. Indeed many existing experiments could be considered as first steps towards classical realizations/emulations of quantum measurements Dembowski et al. (2001, 2004); Doppler et al. (2016). The role of time could be played by a space coordinate along which the non-Hermitian term can be modulated as done in photonic systems Rechtsman et al. (2013).
Final remarks.– We have discussed a framework for a dynamical model of a quantum measurement where collapse emerges from an evolution dictated by the Schrödinger equation. A quantum measurement is seen as resulting from the interaction with other degrees of freedom introducing a non-Hermitian term to an otherwise Hermitian Hamiltonian. The nature of this interaction, having a finite lifespan, motivates using a time-dependent Hamiltonian which, throughout the interaction, describes a closed loop in parameter space. As a possible mechanism we signal the phenomenon called chiral state conversion, which is enabled by exceptional points and leads to the required destruction of superposition states. For a $2\times 2$ system, the outcome can be switched by changing the encircling of the loop from clockwise to anticlockwise. The probability of taking one option or the other, which amounts to choosing how the action of the meter takes place, needs to be given by the initial state so as to recover the usual predictions of quantum mechanics. This makes the dynamic equation also non-linear, which is consistent with other studies Mertens et al. (2021).
Chiral mode conversion has been extensively studied for $2\times 2$ systems, but there is still much to understand for systems with more degrees of freedom. Another mechanism that may come handy is the non-Hermitian skin effect Martinez Alvarez et al. (2018); Yao and Wang (2018), which essentially leads to an anomalous localization at one edge of a pristine lattice. In some cases one might tune the non-Hermitian term to select a few or even a single eigenstate. There has been a surge of interest in this phenomenon, which has been shown to have a topological origin Okuma et al. (2020); Zhang et al. (2020), see for example Okuma and Sato (2023); Bergholtz et al. (2021); Lin et al. (2023) and references therein.
We hope that our discussion may trigger a fresh look to this enduring enigma. Perhaps the challenge ahead is to take the measurement out of the quantum realm and down to an experiment with classical objects.
Acknowledgements.
L.E.F.F.T. thanks Joaquín Márquez Delgado, Diego Bautista Avilés, Igor Gornyi, Yuval Gefen, and Mikhail Kiselev for useful discussions, and C. Cormick for useful comments. L.E.F.F.T. acknowledges financial support by ANID FONDECYT (Chile) through grant 1211038, The Abdus Salam International Center for Theoretical Physics and the Simons Foundation, and by the EU Horizon 2020 research and innovation program under the Marie-Sklodowska-Curie Grant Agreement No. 873028 (HYDROTRONICS Project). ICN2 is funded by the CERCA programme / Generalitat de Catalunya, and is supported by the Severo Ochoa Centres of Excellence programme, Grant CEX2021-001214-S, funded by MCIN / AEI / 10.13039.501100011033. This work is also supported by MICIN with European funds‐NextGenerationEU (PRTR‐C17.I1) and by 2021 SGR 00997, funded by Generalitat de Catalunya.